OpenStudy (ksaimouli):

solve for x where 0<=x<360. your answer should be correct to teh nearest tenth of a degree a.3cosx+9=7

OpenStudy (anonymous):

cos x = -2/3

OpenStudy (ksaimouli):

yes i got 131.8

OpenStudy (anonymous):

I have seen this before .... begin by subtracting 9 from both sides what is next?

OpenStudy (ksaimouli):

but i want second value

OpenStudy (anonymous):

131.8 and 228.2

OpenStudy (ksaimouli):

how to get 2nd one ie 228.2

OpenStudy (asnaseer):

cos(x) = cos(360-x)

OpenStudy (anonymous):

cosine is negative in the 2nd and third quadrants

OpenStudy (ksaimouli):

yes

OpenStudy (anonymous):

2/3 gives 48.2 in first quadrant so third quadrant values is 180 + 48.2 = 228.2

OpenStudy (anonymous):

and 2nd quad is 180-x = 180-48.2 = 131.8

OpenStudy (asnaseer):